OpenStudy (anonymous):

you have 3/4 cups of milk. how much more do you need? 1 1/4 cups of milk.

OpenStudy (anonymous):

is ur q if u have 3/4 cup, how much more do u need to have 1 1/4 cup? is that what ur asking?

OpenStudy (anonymous):

no the 1 1/4 was the other one that u need to add or watever w/ the 3/4

OpenStudy (anonymous):

o so are u asking 3/4 + 1 1/4 ??

OpenStudy (anonymous):

yes, ithink ur sooposed 2 add it..

OpenStudy (anonymous):

ok if thats the case convert 1 1/4 to an improper fraction. 5/4+3/4 = 8/4=2. but would u mind rephrasing the q again

OpenStudy (anonymous):

just to make sure

OpenStudy (anonymous):

rephasing what// sorry

OpenStudy (anonymous):

the q

OpenStudy (anonymous):

oh nvm i got it thx u!!

OpenStudy (anonymous):

yw